# Digital Over-the-Air Federated Learning in Multi-Antenna Systems

### v2

@@ IV-B Robustness of MLP Predictor
IV-C Impacts of Channel Conditions and SNR
We next investigate the impact of channel conditions on the performance of our methodology. In Fig. 9, we show the performance of the proposed method on the CIFAR-10 dataset under various channel models. In particular, we consider the Rayleigh fading channel and the Rician fading channel. For each model, we consider three different settings, varying the noise variance $\sigma^{2}=-90$ dBm, pathloss exponent $\beta$, and Rician factor $k$. Overall, from this figure, we can see that the proposed algorithm is reasonably robust to the specific channel conditions, with the testing accuracy staying within a window of ±$\pm 0.04$. The best performance occurs when the channel is Rician and $k=10$. This is consistent with the Rician factor being the ratio of the power of the dominant path component to the power of the whole signal. Thus, $k=10$ is the case of the largest dominant path, which reduces FL transmission errors and improves the FL performance.
Figure 9: Impact of Rician (dashed lines) and Raleigh (solid lines) channel models on training performance, for the CIFAR-10 dataset under an IID data partition. Overall, we see that our methodology is reasonably robust to variations in the channel model.
@@ IV-C Impacts of Channel Conditions and SNR
IV-D Impact of Network Size
@@ V Conclusion
VI Appendix
VI-A Proof of Theorem 1

### v3

@@ IV-C Robustness of MLP Predictor
IV-D Impacts of Channel Conditions and SNR
We next investigate the impact of channel conditions on the performance of our methodology. In Fig. 9, we show the performance of the proposed method on the Fashion-MNIST dataset under various channel models. In particular, we consider the Rayleigh fading channel and the Rician fading channel. For each model, we consider three different settings, varying the noise variance $\sigma^{2}=-90$ dBm, pathloss exponent $\beta$, and Rician factor $k$. Overall, from this figure, we can see that the proposed algorithm is reasonably robust to the specific channel conditions, with the testing accuracy staying within a window of ±$\pm 0.04$. The best performance occurs when the channel is Rician and $k=10$. This is consistent with the Rician factor being the ratio of the power of the dominant path component to the power of the whole signal. Thus, $k=10$ is the case of the largest dominant path, which reduces FL transmission errors and improves the FL performance.
Figure 9: Impact of Rician (dashed lines) and Raleigh (solid lines) channel models on training performance, for the Fashion-MNIST dataset under an IID data partition. Overall, we see that our methodology is reasonably robust to variations in the channel model.
@@ IV-D Impacts of Channel Conditions and SNR
IV-E Impact of Network Size
@@ V Conclusion
Appendix A Proof of Theorem 1
@@
Appendix B Latency Models
We present here the models employed for calculating latency/delay in the experiments. For our algorithm, the latency encompasses the time of updating and transmitting both the MLP and the local FL models. For the analog FL and BPSK FL baselines, the latency consists of the time of updating and uploading the FL parameters only. Our models of the four components are given below.
MLP updating time: The updating time of the MLP depends on the computational complexity of the MLP and the computational power of each device. The computational complexity of the MLP depends on the size of input $\bm{g}^{\prime}_{t-1}$ and output ${\rm\Delta}{\widetilde{\bm{w}}^{\prime}}_{k,t}$, as well as the number of the neurons in the hidden layer. As discussed in Section III-B, the sizes of $\bm{g}^{\prime}_{t-1}$ and ${\rm\Delta}{\widetilde{\bm{w}}^{\prime}}_{k,t}$ are both $V^{\prime}$, and the number of neurons in the hidden layer is $D$. Then, following the latency models in [63], the updating time of the MLP can be modeled as
$L_{\rm MLP,U}=\frac{2\alpha^{2}V^{\prime}\!D}{\varepsilon f_{\rm D}}\rho,$ (26)
where $2V^{\prime}D$ is the total number of parameters, $f_{\rm D}$ is the CPU clock frequency of a device (in cycles/sec), $\varepsilon$ is the number of parameter update operation executed by one CPU cycle, $\rho$ is the time-consumption coefficient depending on the specific chip of each device, and $\alpha$ is the quantization precision.
FL updating time: The updating time of the FL model similarly depends on the computational complexity of the adopted model and the computational power of each device. For Fashion-MNIST and CIFAR-10, we employ a series of convolutional layers and one linear layer. The training complexity depends on the specific structure of each layer. In each convolutional layer $i$, let $M_{i}$ denote the size of the convolutional kernel, $K_{i}$ denotes the side length of the output feature map generated by each convolutional kernel, and $C^{\rm in}_{i}$ and $C^{\rm out}_{i}$ denote the number of input channels and output channels, respectively. Then, the number of operations required for training one of the models is $O_{\rm FL}=\sum\limits_{i=1}^{I}M^{2}_{i}\times K_{i}^{2}\times C^{\rm in}_{i}\times C^{\rm out}_{i}+C^{\rm out}_{I+1}\times L_{\rm out}$, where $L_{\rm out}$ is the output dimension of the entire network and $I$ is the number of convolutional layers [65]. Then, the updating time of the FL model is given by
$L_{\rm FL,U}=\frac{O_{\rm FL}\alpha^{2}}{\varepsilon f_{\rm D}}\rho.$ (27)
MLP transmission time: The transmission time of the MLP depends on the size of the MLP and the channel states. Given the number of parameters is $2V^{\prime}D$, the transmission time of the MLP can be modeled as
$L_{\rm MLP,T}=\frac{l\left(2\alpha V^{\prime}D\right)}{C},$ (28)
where $l\left(\cdot\right)$ denotes the digital pre-processing function described in Sec. II-A. $C$ is the channel capacity (in bits/sec), which is given for MIMO channels as [64]
$C=\mathbb{E}\left[{\rm log}{\;\rm det}\left({\textbf{I}_{N_{r}}}+\frac{{\rm SNR}}{N_{t}}\right)\textbf{H}\textbf{H}^{*}\right],$ (29)
where ${\rm SNR}=P_{0}/\sigma^{2}$ is the common signal-to-noise ratio (SNR) at each receive antenna with $\sigma^{2}$ being the variance of the additive white Gaussian noise, $N_{t}$ and $N_{r}$ are the numbers of transmitting and receiving antennas, respectively, and $\mathbf{H}$ is the MIMO channel vector.
FL transmission time: The transmission time of the FL model depends on the size of the adopted FL model and the channel state. Given the neural network structures, the transmission delay can be modeled as
$L_{\rm FL,T}=\frac{l\left(V\alpha\right)}{C},$ (30)
where $C$ is as defined above and $V=\sum_{i=1}^{I}M_{i}^{2}\times C_{i}^{\mathrm{in}}\times C_{i}^{\text{out }}+C_{I+1}^{\text{out }}\times L_{\text{out }}$ is the number of parameters in the adopted FL model.
For the CIFAR-10 task, $I=17$, $V=1.17\times 10^{7}$, and $O_{\rm FL}=1.7\times 10^{12}$, and for the Fashion MNIST task, $I=5$, $V=9.2\times 10^{5}$, and $O_{\rm FL}=1.2\times 10^{9}$. The value of $\alpha$ is variable for different baselines: for BPSK, $\alpha=1$ (i.e., single bit precision), for analog FL, $\alpha=32$ (i.e., full precision computation), and for our proposed method using 64 QAM, $\alpha=6$. We set $\varepsilon=10^{5}$, $\rho=1$, and $f_{\rm D}=1$ GHz in our experiments.